40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  SUS Lab eLog  Not logged in ELOG logo
Message ID: 1861     Entry time: Wed Nov 11 12:55:18 2020
Author: Paco 
Type: Lab Infrastructure 
Category: General 
Subject: 1064 pump power curve 

Log of the output power vs current in the 1064 nm (Innolight) pump laser. The crystal temperature was set to 45.5 C, and the current limit is set to 2.1 A

Attachment 1: 2020_11_10_mephisto_test.png  69 kB  | Show | Show all
ELOG V3.1.3-